The typical concentration of Mg2+ in blood is 3 mEq/L.
How many milligrams of Mg2+ are in 380 mL of blood?
Answer #1

3 mEq / litre

volume = 380 ml or 0.380 litre

mEq = 3*0.380

= 1.14mEq

or 1.14*10^-3 Eq

now,

for Mg2+ = 24 / 2

= 12 g

mass = 1.14*10^-3 * 12

= 0.01368 g

= 13.68 mg

any query please comment

if satisfied please rate it thanks